# Neural Priming for Sample-Efficient Adaptation

### v1

@@ Abstract
We propose ${\rm Neural~{}Priming}$, a technique for adapting large pretrained models to distribution shifts and downstream tasks given few or no labeled examples. Presented with class names or unlabeled test samples, Neural Priming enables the model to recall and conditions its parameters on relevant data seen throughout pretraining, thereby priming it for the test distribution. Neural Priming can be performed at test time in even for pretraining datasets as large as LAION-2B. Performing lightweight updates on the recalled data significantly improves accuracy across a variety of distribution shift and transfer learning benchmarks. Concretely, in the zero-shot setting, we see a $2.45\%$ improvement in accuracy on ImageNet and $3.81\%$ accuracy improvement on average across standard transfer learning benchmarks. Further, using our test time inference scheme, we see a $1.41\%$ accuracy improvement on ImageNetV2. These results demonstrate the effectiveness of Neural Priming in addressing the common challenge of limited labeled data and changing distributions. Code is available at github.com/RAIVNLab/neural-priming.
@@ 1 Introduction
${\rm Neural~{}Priming}$ is flexible and can be used with variable degrees of information about the downstream distribution. When the model has language-only task descriptions, our approach can efficiently retrieve a priming pool of relevant examples from the pre-training set and attune the model to this data. At inference time, given a set of test images to classify, Neural Priming is able to further filter the priming pool down using these examples specific to the test distribution. When we have access to labeled examples in the few-shot setting, ${\rm Neural~{}Priming}$ can further filter the priming pool through agreement between our language-only retrieval and those of our ground-truth labeled examples.
@@
  • ${\rm Neural~{}Priming}$ also enables transductive learning and improves performance on standard distribution shift datasets by 2.51% on average, all without using any additional data (Section 4.3).
@@ 2.2 Distribution Shifts
The most significant recent improvement in effective robustness [41] is the introduction of open-vocabulary models. At its time of release, CLIP [40] achieved unprecedented effective robustness on a variety of distribution shifts. Studies have suggested that these models achieve high effective robustness through their data distribution [9], a result of training on large amounts of web-scraped data. However, these models are still worse at downstream tasks than models fine-tuned on in-distribution data. Moreover, fine-tuning on downstream data causes robustness on other data distributions to deteriorate [40, 50]. Many mitigation methods have been proposed to such as Wise-FT, FLYP, LP-FT, and model surgery [50, 15, 28, 29]. Our paper differs from these methods in goal: whereas they seek to keep model robustness while gaining the benefits of fine-tuning on task-specific data, we seek to gain the benefits of fine-tuning while not collecting any in-distribution data. Hence these methods are complementary to Neural Priming, and indeed we employ Wise-FT in our model attunement procedure.
@@ 2.5 Retrieval-Augmented Models
In language, works have demonstrated the effectiveness of retrieval from text corpuses or structured data for tasks such as question answering [3, 16, 25]. In general, these methods seek to recover facts either from a large corpus or knowledge graph, then use those to complete tasks. This differs from our scenario, where exact examples at inference time do not necessarily exist in the pre-training corpus. REACT is a retrieval-augmented method for open-vocabulary models which also attempts to learn fine-tune  [30]. We differ from REACT in a few key aspects. Firstly, we do not add parameters to our CLIP model, which remains frozen during our “attunement" procedure (described in Section 3.2). Secondly, our approach is more efficient, both in computationally and in terms of number of samples, allowing us to use it for additional improvements at test time (Section 3.1.2).
@@ 3 Method
${\rm Neural~{}Priming}$ is the process of retrieving relevant information from the pre-training dataset and leveraging it for a specific task. We study it in the context of vision-language contrastive pre-training, so the form our task description takes is a set of class names, $\mathcal{C}$, already in natural language. A CLIP model [40] comprises of a vision embedding model, $V$, and a language embedding model, $L$, each producing a vector representation in $\mathcal{R}^{d}$. The pre-training dataset, $\mathcal{D}$, consists of a large number of image-text pairs collected from the web. The text component can be noisy, potentially containing irrelevant or inaccurate information about the image content.
@@ 3.1.1 Leveraging Natural Language Task Information
The goal of this step is to collect an initial pool of images relevant to the task at hand given the previously defined natural language description $\mathcal{C}$. For example, if our task is a set of dog breeds, ideally we would collect sets of images belonging to those breeds and label them accordingly. A simple way to prime is by using retrieval to gather relevant data points from our pre-training dataset. An existing method for language-based retrieval involves using the CLIP text embedding of a class description $c\in\mathcal{C}$ for retrieval using semantic similarity scores on the pre-training set [2]. However, with neural priming, prioritizing precision over recall is crucial, considering the size, diversity, and noise of the pre-training dataset. This form of semantic retrieval has a major downside: it is not clear where to threshold similarity scores to retrieve the most relevant images. Threshold too late and we allow unrelated images to be included in our pool. Further, this threshold is often specific to a category, making it infeasible to search at scale.
Our approach to language-based priming is to search for the existence of the class name, $c\in\mathcal{C}$, in the captions of our pre-training dataset to retrieve images relevant to a particular category. We organize these image-text pairs into separate categorical clusters $\{B_{c}\}$ according to the class name $c$ mentioned in their captions. This approach has a few advantages over semantic retrieval: 1. After setting up an inverted index search structure for text retrieval [26, 45], exact string matching is far faster than semantic retrieval, even when approximate nearest neighbor strategies are employed [24], 2. with exact string search, the category boundary is clear and therefore does not require per category tuning, 3. the retrieval results are overall qualitatively more relevant. Finally, to take advantage of the semantic understanding of our CLIP model, we then hard filter the priming pool. We do this by constructing a “zero-shot” CLIP classifier, as defined by Radford et al. [40], and removing examples from categorical clusters that do not align with their label according to the CLIP model.
@@ 3.1.2 Leveraging Image Information at Test Time
At inference time, the model can narrow the relevant priming pool even further by utilizing information about the test distribution. To do this, given an image $x$ in our test set, we compute the cosine similarity using our CLIP image encoder $V$, $\cos(V(x),V(y))$ for every $y\in P$, our priming pool. We retrieve examples with the top-$k$ cosine similarity scores ($k=10$ in most of our experiments). We do this for collectively for every image in the test set and collect the retrievals to form a filtered priming pool. If an example is retrieved twice, we de-duplicate them in the final priming pool. Since we do this for all images in the test set, we consider this the transductive setting to align with prior work [13, 6].
@@ 3.2 Attuning CLIP to the Priming Pool
The goal of this step is to modify our CLIP model to take advantage of the data in the priming pool $P$. Note that we do not add extra parameters to the final CLIP classifier, unlike [30], and the backbone weights remain unchanged.
We first construct the task-specific zero-shot linear head $W_{z}\in\mathbb{R}^{d\times n}$, where $d$ is the feature dimension and $n$ is the number of classes, for our vision encoder using the text model as in [40]. To get logits for a particular example, $x$, we compute $W_{z}\cdot L(x)$, so our prediction is $\operatorname*{arg\,max}_{c}W_{z}\cdot L(x)$.
To attune our CLIP model to the priming pool, we perform nearest-class mean (NCM) [32] on all retrieved examples per category to get a pool specific classification head. Namely for a given class $c$, we compute a centroid $\tilde{y}_{c}={1\over|B_{c}|}\sum_{x\in B_{c}}L(x)$ and then normalize this centroid to produce a class embedding $y_{c}=\tilde{y}_{c}/\|y_{c}\|$. We define the collection of centroids as matrix $W_{ft}=[y_{c}]_{c}\in\mathbb{R}^{d\times n}$. To expand this to few-shot scenarios, we mix the labeled data into the corresponding categorical clusters before performing NCM. Finally, we mix $W_{z}$ and $W_{ft}$ using a mixing coefficient $\alpha\in[0,1]$ as $W_{\alpha}=(1-\alpha)\cdot W_{ft}+\alpha\cdot W_{z}$, which is our final classification head. We choose alpha according to an exponential scaling heuristic $\alpha=e^{-|P|^{2}/\sigma}$. We discuss our choice of this heuristic in the appendix. Intuitively, if we do not have much data in our priming pool, we want it to influence our model less. Using $\alpha$ mixing when fine-tuning CLIP models has been shown to reduce the effects of distribution shift [50].
@@ 4.1 Datasets and Architectures
We perform our experiments with OpenCLIP models [50] trained on LAION-2B and 400M [43]. We choose OpenCLIP because their pretrain datasets are publicly available, therefore we can ensure no additional data is introduced. The model architecture reported in the main paper is the B-16 variant trained on LAION-2B unless otherwise stated and we report L-14 and B-32 in the Appendix C.
@@ 4.2 Zero-shot Results
We also compare with VLM [33] and CuPL [39], two zero-shot prompt-tuning methods which obtain natural language descriptions of each class using language models, and a retrieval with fine-tuning baseline. For implementation details on how the retrieval and fine-tuning is performed see Appendix H. Interestingly, we find that Neural Priming is complementary to existing prompt-tuning methods. The accuracy improvements from CuPL and VLM are additive with Neural Priming. For example, CuPL and Neural Priming each improve performance by 3.78% and 7.17% respectively on FGVCAircraft. Ensembling the methods results in 10.74% improvement over the baseline (Table 2). This surprising result suggests that the textual class descriptions in CuPL and VLM provide unique information to the model that differs from the information obtained from the images in the priming set.
@@
To be precise, when we refer to term “shot number” throughout the experiments section, we mean the number of labeled examples from the target training set. We do not consider images retrieved from LAION-2B as shots in this setting because they can be obtained from the pre-training set automatically.

### v3

@@ Abstract
We propose ${\rm Neural~{}Priming}$, a technique for adapting large pretrained models to distribution shifts and downstream tasks given few or no labeled examples. Presented with class names or unlabeled test samples, Neural Priming enables the model to recall and conditions its parameters on relevant data seen throughout pretraining, thereby priming it for the test distribution. Neural Priming can be performed at inference, even for pretraining datasets as large as LAION-2B. Performing lightweight updates on the recalled data significantly improves accuracy across a variety of distribution shift and transfer learning benchmarks. Concretely, in the zero-shot setting, we see a $2.45\%$ improvement in accuracy on ImageNet and $3.81\%$ accuracy improvement on average across standard transfer learning benchmarks. Further, using Neural Priming at inference to adapt to distribution shift, we see a $1.41\%$ accuracy improvement on ImageNetV2. These results demonstrate the effectiveness of Neural Priming in addressing the challenge of limited labeled data and changing distributions. Code is available at https://github.com/RAIVNLab/neural-priming.
@@ 1 Introduction
${\rm Neural~{}Priming}$ is flexible and can be used with variable degrees of information about the downstream distribution. When the model has language-only task descriptions, our approach can efficiently retrieve a priming pool of relevant examples from the pre-training set and attune the model to this data. At inference time, given a set of test images to classify, Neural Priming is able to use test examples to adapt the priming pool to distribution shifts. When we have access to training examples in the few-shot setting, ${\rm Neural~{}Priming}$ can filter the priming pool to align with the training distribution.
@@
  • ${\rm Neural~{}Priming}$ also enables transductive learning and improves performance on standard distribution shift datasets by $2.51\%$ on average, all without using any additional data (Section 4.3).
@@ 2.2 Distribution Shifts
The most significant recent improvement in effective robustness [41] is the introduction of open-vocabulary models. At its time of release, CLIP [40] achieved unprecedented effective robustness on a variety of distribution shifts. Studies have suggested that these models achieve high effective robustness through their data distribution [9], a result of training on large amounts of web-scraped data. However, these models are still worse at downstream tasks than models fine-tuned on in-distribution data. Moreover, fine-tuning on downstream data causes robustness on other data distributions to deteriorate [40, 52]. Many mitigation methods have been proposed to such as Wise-FT, FLYP, LP-FT, and model surgery [52, 15, 28, 29]. Our paper differs from these methods in goal: whereas they seek to keep model robustness while gaining the benefits of fine-tuning on task-specific data, we seek the benefits of fine-tuning while not collecting any in-distribution data. Hence these methods are complementary to Neural Priming, and we employ Wise-FT in our model attunement procedure.
@@ 2.5 Retrieval-Augmented Models
In language, works have demonstrated the effectiveness of retrieval from text corpora or structured data for tasks such as question answering [3, 16, 25]. In general, these methods seek to recover facts either from a large corpus or knowledge graph, then use those to complete tasks. This differs from our scenario, where exact examples at inference time do not necessarily exist in the pre-training corpus. REACT [30] and SuS-X [50] are retrieval-augmented methods for open-vocabulary models which use search to fine-tune with relevant examples [30]. We differ from Liu et al. [30] in that they add a substantial number of new parameters whereas we do not. Additionally, our approach is significantly more efficient, both computationally and in terms of number of samples, enabling use at inference for additional improvement (Section 3.1.2). We differ from [50] in that their work uses semantic retrieval whereas ${\rm Neural~{}Priming}$ leverages language for fast initial filtering and image search for accurate retrieval. Further, Neural Priming shows that models can improve by revisiting examples seen throughout pretraining whereas other works retrieve new examples from external datasets.
@@ 3 Method
${\rm Neural~{}Priming}$ is the process of retrieving relevant information from the pre-training dataset and leveraging it for a specific task. We study it in the context of vision-language contrastive pre-training, so the form our task description takes is a set of class names, $\mathcal{C}$, already in natural language. A CLIP model [40] consists of a vision embedding model, $V$, and a language embedding model, $L$, each producing a vector representation in $\mathbb{R}^{d}$. The pre-training dataset, $\mathcal{D}$, consists of a large number of image-text pairs collected from the web. The text component can be noisy, potentially containing irrelevant or inaccurate information about the image content.
@@ 3.1.1 Leveraging Natural Language Task Information
The goal of this step is to collect an initial pool of images relevant to the task at hand given the previously defined natural language description $\mathcal{C}$. For example, if our task is a set of dog breeds, ideally we would collect sets of images belonging to those breeds and label them accordingly. A simple way to prime is by using retrieval to gather relevant data points from our pre-training dataset. An existing method for language-based retrieval involves using the CLIP text embedding of a class description $c\in\mathcal{C}$ for retrieval using semantic similarity scores on the pre-training set [2, 42]. However, with neural priming, prioritizing precision over recall is crucial, considering the size, diversity, and noise of the pre-training dataset. This form of semantic retrieval has a major downside: it is not clear where to threshold similarity scores to retrieve the most relevant images. Threshold too late and we allow unrelated images to be included in our pool. Further, this threshold is often specific to a category, making it infeasible to search at scale.
Our approach to language-based priming is to search for the existence of the class name, $c\in\mathcal{C}$, in the captions of our pre-training dataset to retrieve images relevant to a particular category. We organize these image-text pairs into separate categorical clusters $\{B_{c}\}$ according to the class name $c$ mentioned in their captions. This approach has a few advantages over semantic retrieval: 1. After setting up an inverted index search structure for text retrieval [26, 46], exact string matching is far faster than semantic retrieval, even when approximate nearest neighbor strategies are employed [24], 2. with exact string search, the category boundary is clear and therefore does not require per category tuning, 3. the retrieval results are overall qualitatively more relevant. Finally, to leverage the semantic understanding of our CLIP model, we filter the priming pool using CLIP similarity score. We do this by constructing a “zero-shot” CLIP classifier, as defined by Radford et al. [40], and removing examples from categorical clusters that do not align with their label according to the CLIP model.
@@ 3.1.2 Leveraging Image Information at Test Time
At inference time, the model can narrow the relevant priming pool even further by utilizing information about the test distribution. To do this, given an image $x$ in our test set, we compute the cosine similarity using our CLIP image encoder $V$, $\cos(V(x),V(y))$ for every $y\in P$, our priming pool. We retrieve examples with the top-$k$ cosine similarity scores ($k=10$ in most of our experiments). We do this collectively for every image in the test set and collect the retrievals to form a filtered priming pool. If an example is retrieved twice, we de-duplicate them in the final priming pool. Since we do this for all images in the test set, we consider this the transductive setting to align with prior work [13, 6].
@@ 3.2 Attuning CLIP to the Priming Pool
The goal of this step is to modify our CLIP model to take advantage of the data in the priming pool $P$. We first construct the task-specific zero-shot linear head $W_{z}\in\mathbb{R}^{d\times n}$ using the text encoder and the natural language names of each class, where $d$ is the feature dimension and $n$ is the number of classes. To get logits for a particular example, $x$, we compute $W_{z}\cdot L(x)$, so our prediction is $\operatorname*{arg\,max}_{c}W_{z}\cdot L(x)$.
To attune our CLIP model to the priming pool, we perform nearest-class mean (NCM) [32] on all retrieved examples to obtain a classification head from the priming pool. Namely for a given class $c$, we compute a centroid $\tilde{y}_{c}={1\over|B_{c}|}\sum_{x\in B_{c}}L(x)$ and then normalize this centroid to produce a class embedding $y_{c}=\tilde{y}_{c}/\|y_{c}\|$. We define the collection of centroids as matrix $W_{ft}=[y_{c}]_{c}\in\mathbb{R}^{d\times n}$. To expand this to few-shot scenarios, we mix the labeled data into the corresponding categorical clusters before performing NCM. Finally, we ensemble $W_{z}$ and $W_{ft}$ using a mixing coefficient $\alpha\in[0,1]$ as $W_{\alpha}=(1-\alpha)\cdot W_{ft}+\alpha\cdot W_{z}$, which is our final classification head. We choose alpha according to a heuristic $\alpha=e^{-|P|/\sigma^{2}}$ which can be derived from a Bayesian prior over the text features. We experiment with varying values of $\sigma$ in Table 10. We also find that $\alpha$ can be effectively chosen through cross-validation on a held-out portion of the retrieval set. Intuitively, if we do not have much data in our priming pool, we want it to influence our model less. We use NCM as the classifier as it has shown to be sample-efficient [47]. For comparison to a WISE-FT classifier  [52] see Table 3.
@@ 4.1 Datasets and Architectures
We perform our experiments with OpenCLIP models [52] trained on LAION-2B and 400M [44]. We choose OpenCLIP because their pretrain datasets are publicly available, therefore we can control what data is introduced to the model. The model architecture reported in the main paper is the B-16 variant trained on LAION-2B unless otherwise stated and we report L-14 and B-32 in the Appendix C.
@@ 4.2 Zero-shot Results
We also compare with VLM [33] and CuPL [39], two zero-shot prompt-tuning methods which obtain natural language descriptions of each class using language models, and a retrieval with fine-tuning baseline. For implementation details on how the retrieval and fine-tuning are performed see Appendix H. Interestingly, we find that Neural Priming is complementary to existing prompt-tuning methods. The accuracy improvements from CuPL and VLM are additive with Neural Priming. For example, CuPL and Neural Priming each improve performance by 3.78% and 7.17% respectively on FGVCAircraft. Ensembling the methods results in 10.74% improvement over the baseline (Table 2). This surprising result suggests that the textual class descriptions in CuPL and VLM provide unique information to the model that differ from the information obtained from the images in the priming set.
@@
To be precise, when we refer to term “shot number” throughout the experiments section, we mean the number of labeled examples from the target training set. We do not consider images retrieved from LAION-2B as shots in this setting because they are obtained from the pre-training set.
@@ Appendix A Implementation Details
Table 3: Comparing Neural Priming to other classification approaches. NCM uses only the image features. WISE-FT fine-tunes a linear classifier and ensembles it with the text classifier.
Shots Neural Priming NCM Wise-FT
0 71.38 68.3 67.75
1 71.56 68.89 68.14
2 71.81 68.85 68.25
5 72.03 69.12 68.78
10 72.36 70.12 69.02
@@ Appendix H Other Implementation Details
$\sigma$ ImageNet Stanford Cars FGVC-Aircraft Flowers102
1 68.9 87.54 25.9 71.94
10 69.1 88.98 29.75 74.97
100 70.75 89.3 33.03 79.81
1000 70.64 89.11 31.32 78.81
Table 10: An ablation over the sigma value used to calculate the mixing coefficient between the image and text features. A larger sigma value is associated with a stronger prior for the text features.